# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

### v1

@@ 4 Discussion
  • Some TDEs (e.g., AT2018hco, the second radio flare in AT2019dsg) have radio outflow launch times that roughly coincide with MOSFiT’s predicted transition to a highly sub-Eddington state $\dot{M}<0.03\dot{M}_{Edd}$. However, we find that other delayed radio outflows are launched up to several hundred days later than the MOSFiT predictions for the timing of this state transition (e.g., in ASASSN-14ae).
  • Similarly, while the radio launch times of a few TDEs in our sample may be marginally consistent with MOSFiT’s predicted super-Eddington phase $\dot{M}>0.3\dot{M}_{Edd}$ (ASASSN-15oi’s first flare, iPTF16fnl, PS16dtm, AT2019ezh), the radio emission in most of our delayed radio TDEs can only be explained as outflows powered by super-Eddington accretion if the super-Eddington phase is delayed by several hundred days relative to the MOSFiT prediction.
@@
  • Many TDEs, including all TDEs in our sample with prompt radio emission and some TDEs with delayed radio emission, are broadly consistent with the Fundamental Plane of black hole activity.
@@
Broadly speaking, there are two classes of models that have been proposed to explain radio emission in TDEs: models in which the radio emission is directly powered by accretion-driven outflows (e.g. Strubbe & Quataert 2009; Giannios & Metzger 2011; De Colle et al. 2012; Alexander et al. 2016; van Velzen et al. 2016; Metzger 2022), and those where it originates from some other aspect of the disruption process unrelated to SMBH accretion. The latter category includes emission from shocks between the unbound portion of the stellar debris stream and the ambient environment (Krolik et al., 2016), emission from outflows launched by stream-stream collisions during the circularization of the bound stellar debris (Lu & Bonnerot, 2020), and sudden brightenings of a shock that encounters density variations in the circumnuclear environment (Zhuang et al., 2024). In the first category, we might expect correlations between the radio behavior and other probes of the accretion rate (such as the optical or X-ray emission), while in the second category we would not expect such correlations. As we explore in this section, the existence of the various multi-wavelength connections uncovered in our analysis thus suggest that the late-time radio properties of our sample are driven by accretion onto the SMBH, rather than by environmental interactions.
4.1 Overview of accretion-driven jets and outflows in TDEs
@@
As discussed in Giannios & Metzger (2011), there are two phases after a TDE where an accretion-driven relativistic jet or other radio-producing outflow may be formed. Firstly, if the process of debris circularization and disk formation is sufficiently rapid, then a jet may accompany the early super-Eddington phase ($\dot{M}\gg\dot{M}_{\rm Edd})$, in analogy with gamma-ray bursts. Such an early, super-Eddington jetted phase is likely responsible for highly energetic jetted TDEs such as Swift J1644+57 (De Colle et al., 2012). This early jet phase ends once the disk becomes geometrically thin at $\dot{M}\lesssim 0.3\dot{M}_{\rm Edd}$ and can no longer retain its magnetic flux (e.g., McClintock et al. 2006; Tchekhovskoy et al. 2014). This transition typically occurs hundreds of days post-disruption, consistent with the observed shut-off time of the X-rays in jetted TDEs (Zauderer et al., 2013; Cenko et al., 2012; Eftekhari et al., 2024). Super-Eddington accretion has also been proposed to power non-relativistic winds in TDEs that produce low-luminosity radio emission at early times (e.g., Strubbe & Quataert 2009; Alexander et al. 2016). The presence of a wind may explain why only powerful super-Eddington jets have been seen at early times: if the star’s orbital plane is initially misaligned with the black hole spin axis, the jet will precess and less powerful jets may only be able to break out of the wind after the disk undergoes sufficient Lense-Thirring precession to come into alignment with the spin (Teboul & Metzger, 2023; Lu et al., 2024). This will then delay any radio emission from less powerful super-Eddington jets. Radio emission from powerful relativistic jets may also appear delayed if the jets are viewed off-axis, as relativistic beaming will suppress the emission at early times.
@@
Alternatively, a jet may form at very late times, once the accretion rate decreases to $\dot{M}\lesssim 0.03\dot{M}_{\rm Edd}$ (Giannios & Metzger, 2011). When the accretion rate becomes this low, the accretion flow is predicted to transition back to radiatively inefficient and geometrically thick, which can again power outflows (e.g., Narayan & Yi 1995; Blandford & Begelman 1999). Such behavior is supported by detections of radio jets from X-ray binaries (XRBs) in the low-hard state and bright radio emission consistent with compact jets from AGN at low Eddington ratios (e.g., Fender et al. 1999; Falcke et al. 2004; Sikora et al. 2007; Corbel et al. 2013). The rapid X-ray and optical spectral and photometric variability observed in changing-look AGN (Matt et al., 2003; Ricci & Trakhtenbrot, 2023) also supports the existence of this state change, as this variability is consistent with the sudden appearance or disappearance of a jet and changing-look AGN are predominantly observed to have Eddington ratios of a few percent, right at this expected transition. (This is much lower than the typical Eddington ratios seen in the general AGN population; MacLeod et al. 2019; Green et al. 2022; Guo et al. 2024.) A delayed radio jet was recently discovered in a changing-look AGN $\sim 6$ years after its original optical outburst, providing direct evidence that these events can power outflows (Meyer et al., 2025). In TDEs, this late jet phase is expected to occur on timescales of years to decades.
@@
We expect our TDEs to adhere to the Fundamental Plane of black hole activity (Section 3.4) during periods where their SMBHs are accreting at a low fraction of the Eddington rate (Falcke et al., 2004; Jonker et al., 2020). Conversely, if the accretion rate is still high during some observed delayed radio brightenings, then we might expect the radio and X-ray emission to deviate from the Fundamental Plane for these TDEs. We may also expect to see a given TDE evolve along the Plane or towards/away from it as $\dot{M}$ changes and the accretion flow possibly undergoes state changes. The radio and X-ray emission of some TDEs also may not follow the Fundamental Plane at any epoch if the radio emission is not directly driven by accretion onto the SMBH (e.g., if prompt emission is produced instead by the shock between the unbound portion of the stellar debris stream and the ISM; Krolik et al. 2016, and/or if late-time emission is due to a previously-launched outflow encountering a change in the ambient environment; Zhuang et al. 2024). In the case of radio emission dominated by a precessing jet whose escape is delayed due to an initial misalignment between the disk and the SMBH spin, we also do not expect any clear correlation between the simultaneous radio and X-ray emission, as the radio onset time is determined by the alignment timescale rather than the accretion rate in these models (Teboul & Metzger, 2023; Lu et al., 2024).
4.2 Origin(s) of the delayed radio emission
@@
Our delayed radio TDEs exhibit radio emission on a wide range of timescales (hundreds to thousands of days post-discovery) and with a wide range of luminosities, suggesting that a single mechanism may not be sufficient to explain the entire sample. In this section, we discuss the various types of accretion-driven outflows that may contribute to the delayed radio emission. As described in Section 3, we constrain the accretion rate as a function of time for the TDEs in our sample via two methods: modeling the optical/UV light curves using MOSFiT, and comparing their X-ray and radio properties directly to the Fundamental Plane of black hole activity.
4.2.1 (Delayed) super-Eddington accretion
@@
We first consider scenarios in which the delayed radio emission observed in our TDE sample is powered by non-relativistic outflows launched during a phase of super-Eddington accretion. While this is unlikely to explain TDEs with radio emission delayed by thousands of days (e.g. ASASSN-14ae) [^14] It is possible to delay peak accretion onto the SMBH by this much under certain assumptions (for example in the models for low-mass SMBHs explored by Guillochon & Ramirez-Ruiz 2015, where the debris stream may only self-intersect after dozens of pericenter passes and debris circularization can therefore be delayed by years). However, delaying the stream-stream collisions and debris circularization phases to this extent also spreads out the resulting mass fallback such that the peak accretion rate is likely no longer super-Eddington, and the optical evolution is likely much slower than observed in ASASSN-14ae., it may be a plausible explanation for TDEs where a radio outflow is launched hundreds days after discovery.
@@
Many of our delayed radio TDEs have larger optical photospheric radii at early times and delayed or absent He lines. Previous work by Nicholl et al. (2022) found that TDE-H may arise from either less complete stellar disruptions (i.e., those with a smaller impact parameter, $b$) or from those in which an accretion disk does not form promptly. Our observed correlation between delayed radio emission and a lack of He emission could thus suggest that less complete stellar disruptions are also more likely to produce delayed radio emission, but we disfavor this interpretation as we find no statistically-significant difference in the $b$ values of our Prompt and Delayed TDE sub-samples (Appendix C). Instead, the fact that our delayed radio subsample both lacks prompt He emission and has larger optical photospheres favors the delayed accretion hypothesis. Nicholl et al. (2022) also found that their TDE-H exhibit larger photospheric radii than other types of TDEs, consistent with previous studies (Hinkle et al., 2020; van Velzen et al., 2021; Charalampopoulos et al., 2022). Nicholl et al. (2022) suggest that this could imply the early optical emission in TDE-H could be powered by collision-induced outflows rather than promptly formed accretion disks, such that accretion onto the SMHB is delayed relative to peak optical light. This would also naturally delay the onset of radio emission produced by accretion-powered outflows.
@@
Alternately, the stellar debris may circularize rapidly without immediately accreting. Instead, the circularized debris may initially form a hot quasi-spherical pressure-supported envelope (Steinberg & Stone, 2024; Price et al., 2024), delaying rapid accretion onto the SMBH until the envelope cools and contracts to form a disk (Loeb & Ulmer, 1997; Coughlin & Begelman, 2014; Metzger, 2022). If the pressure-supported envelope model is correct, the timing of the super-Eddington accretion phase may be delayed by several hundred days relative to the MOSFiT models presented in Section 3.2, suggesting that it may still be possible for super-Eddington accretion to power the delayed radio emission observed in some TDEs even if the radio onset time occurs after MOSFiT’s nominal $\dot{M}\sim 0.3\;\dot{M_{\rm Edd}}$ transition point rather than prior to it (Metzger, 2022; Sarin & Metzger, 2024). In this case, the peak of the X-ray light curve will also be delayed relative to peak optical light and may provide a better estimate of the peak accretion time (Metzger, 2022). This is a likely explanation for the delayed X-rays seen in ASASSN-15oi (Figure 2 and Hajela et al. 2024). In scenarios where the early TDE UV/optical emission instead originates from a quasi-spherical envelope (Metzger, 2022), the initial size of the envelope (and hence its photosphere radius) increases with the SMBH mass; however, the cooling timescale over which the envelope contracts is also longer for lower SMBH mass, such that in the simplest scenario of passive cooling, the photosphere radius at a fixed time after the disruption is roughly independent of the SMBH mass. Given the low significance of our possible correlation between SMBH mass and delayed radio emission, this scenario is also consistent with our results.
4.2.2 Relativistic jets
@@
The rapid radio brightening seen in these events may be powered by a jet (or wind) launched during a period of high accretion onto their SMBHs, when they are not expected to be on the Fundamental Plane. This is broadly consistent with the MOSFiT accretion rate predictions at the outflow launch times for ASASSN-15oi, PS16dtm, and AT2019ehz shown in Figures 2 and 3, although the systematic uncertainties on the SMBH masses for all three TDEs are large. The launch time for AT2018hyz shown in Figure 2 is taken from the non-relativistic outflow modeling by Cendes et al. (2022b) and does not coincide with MOSFiT’s prediction for super-Eddington accretion. This may either suggest that MOSFiT places the time of peak accretion too early, or that AT2018hyz’s radio emission is powered by an off-axis jet launched much earlier than the Cendes et al. (2022b) prediction but not visible until late times because of relativistic beaming (Matsumoto & Piran, 2023) or delayed disk-black hole spin alignment (Teboul & Metzger, 2023; Lu et al., 2024). Further radio monitoring of AT2018hyz, including Very Long Baseline Interferometry observations, may distinguish among these possibilities.
@@
4.2.3 Highly sub-Eddington accretion
@@
Some delayed radio emission in TDEs may also be powered by highly sub-Eddington accretion. To assess this possibility, we must accurately determine the timing of the $\dot{M}\;\sim 0.03\dot{M}_{\rm Edd}$ state transition, when the disk is expected to transition from geometrically thin back to geometrically thick. In our Case 1 prompt accretion scenario, we can straightforwardly compute this timescale using our $\dot{M}$ curves from MOSFiT (Section 3.3). In Case 2 delayed accretion scenarios, the situation is less clear cut. If the cooling envelope model developed by Metzger (2022) is correct, then assuming that the envelope contracts to a disk faster than the mass fall-back reaches $\dot{M}\;\sim 0.03\dot{M}_{\rm Edd}$, and neglecting any spreading disk phase, the timescale estimates for this state transition that we obtain with MOSFiT while neglecting the envelope phase may still be accurate. [^16] We note that even if the MOSFiT prediction for the timing of the $\dot{M}\;\sim 0.03\dot{M}_{Edd}$ transition is underestimated, parameters such as $R_{\rm ph,0}$ that are primarily determined by the photometry around peak optical light are still reliable. We find that some TDEs in our sample (e.g. AT2018hco) do indeed exhibit delayed radio emission consistent with an outflow launched at approximately the same time that MOSFiT predicts this state transition will occur (Figure 2).
@@
However, the existence of UV plateaus in many TDEs that remain well-fit by a thin disk model for years post-disruption (e.g., van Velzen et al. 2019a; Mummery & Balbus 2020; Mummery et al. 2024a, b) may suggest that in reality the $\dot{M}\;\sim 0.03\dot{M}_{Edd}$ transition is also delayed relative to the MOSFiT prediction, at least for some TDEs. [^17] This is not always the case; Mummery et al. (2024a) apply the thin disk model FitTeD to AT2019dsg and find that the onset of its second radio flare approximately coincides with the calculated $\dot{M}\;\sim 0.03\dot{M}_{Edd}$ state transition, just as MOSFiT does (Figure 3). Furthermore, in contrast to the hard X-ray spectra typical of AGN accreting at highly sub-Eddington rates, the X-ray emission from TDEs on timescales of years is often soft, consistent with thermal thin-disk emission extending to late times (e.g., Auchettl et al. 2017; Gezari et al. 2017; Wen et al. 2020; Jonker et al. 2020; Guolo et al. 2023 – but see also Komossa et al. 2004; Saxton et al. 2021 for discussion of X-ray selected TDEs that do show spectral hardening on timescales of years). We therefore must consider the full multiwavelength properties of each TDE in our sample to assess whether MOSFiT can provide an accurate estimate of its state transition timing, or whether other accretion probes (such as the X-ray emission) should be preferred.
@@
The strongest evidence that some of our delayed radio TDEs may have entered the highly sub-Eddington phase comes from comparing their radio and X-ray emission. Many of our delayed radio TDEs (e.g., ASASSN-14ae, AT2018fyk, and AT2019teq) align with the Fundamental Plane at late times (Figure 5). This may indicate that these TDEs have indeed transitioned to an analog of the low-hard state seen in XRBs at late times and, like the AGN and XRBs used to calibrate the original Fundamental Plane relationship, are accreting at a small fraction of the Eddington rate at the time of observation. Indeed, AT2018fyk’s and AT2019teq’s X-ray spectra show significant hardening at late times, consistent with this picture (Wevers et al., 2019b, 2021; Yao & Guolo, 2022; Guolo et al., 2023; Pasham et al., 2024). This is also consistent with the radio non-detections of AT2018fyk at very early times when the X-rays were significantly brighter and softer, as outflows are not expected prior to this state transition when the accretion rate is higher but likely not super-Eddington given the high SMBH mass (Wevers et al., 2019b). ASASSN-14ae is the only TDE in our sample to show a correlated increase in both radio and X-ray emission at late times (Figure 2), but its late-time X-ray emission remains too faint for a detailed spectral analysis (Jonker et al., 2020). ASASSN-14ae is also the oldest TDE in our sample and has the longest delay between peak optical light and first radio detection. As illustrated by Figure 2, in order to explain ASASSN-14ae’s radio emission as an outflow launched by a state transition in the accretion flow at $\dot{M}\sim 0.03\dot{M}_{\rm Edd}$, accretion onto the SMBH must be significantly delayed relative to the $\dot{M}$ curve computed by MOSFiT (by $\sim 2000$ days). Thus, further monitoring of ASASSN-14ae’s radio and X-ray emission is necessary to ascertain the nature of the late radio rise in this TDE.

### v2

@@ IV Discussion
  • Although we are limited by systematics in our SMBH mass measurements, 3-10 of our TDEs (e.g., AT2018hco, the second radio flare in AT2019dsg) have radio outflow launch times that roughly coincide with MOSFiT’s predicted transition to a highly sub-Eddington state $\dot{M}<0.03\,\dot{M}_{\rm Edd}$. However, we find that 3-6 delayed radio outflows are launched up to several hundred days later than the MOSFiT predictions for the timing of this state transition (e.g., in ASASSN-14ae).
  • Similarly, while the radio launch times of a few TDEs in our sample may be marginally consistent with MOSFiT’s predicted super-Eddington phase $\dot{M}>0.3\,\dot{M}_{\rm Edd}$ (ASASSN-15oi’s first flare, iPTF16fnl, PS16dtm, AT2019ezh), the radio emission in most of our delayed radio TDEs can only be explained as outflows powered by super-Eddington accretion if the super-Eddington phase is delayed by several hundred days relative to the MOSFiT prediction.
@@
  • $22$ of our TDEs, including all TDEs in our sample with prompt radio emission and some TDEs with delayed radio emission, are broadly consistent with the Fundamental Plane of black hole activity.
@@
Broadly speaking, there are two classes of models that have been proposed to explain radio emission in TDEs: models in which the radio emission is directly powered by accretion-driven outflows (e.g., Strubbe and Quataert 2009; Giannios and Metzger 2011; De Colle et al. 2012; Alexander et al. 2016; van Velzen et al. 2016; Metzger 2022), and those where it originates from some other aspect of the disruption process unrelated to SMBH accretion. The latter category includes emission from shocks between the unbound portion of the stellar debris stream and the ambient environment (Krolik et al., 2016), emission from outflows launched by stream-stream collisions during the circularization of the bound stellar debris (Lu and Bonnerot, 2020), and sudden brightenings of a shock that encounters density variations in the circumnuclear environment (Zhuang et al., 2025). In the first category, we might expect correlations between the radio behavior and other probes of the accretion rate (such as the optical or X-ray emission), while in the second category we would not expect such correlations. As we explore in this section, the existence of the various multi-wavelength connections uncovered in our analysis thus suggest that the late-time radio properties of our sample are driven by accretion onto the SMBH, rather than by environmental interactions.
IV.1 Overview of accretion-driven jets and outflows in TDEs
@@
As discussed in Giannios and Metzger (2011), there are two phases after a TDE where an accretion-driven relativistic jet or other radio-producing outflow may be formed. Firstly, if the process of debris circularization and disk formation is sufficiently rapid, then a jet may accompany the early super-Eddington phase ($\dot{M}\gg\dot{M}_{\rm Edd})$, in analogy with gamma-ray bursts. Such an early, super-Eddington jetted phase is likely responsible for highly energetic jetted TDEs such as Swift J1644+57 (De Colle et al., 2012). This early jet phase ends once the disk becomes geometrically thin at $\dot{M}\lesssim 0.3\,\dot{M}_{\rm Edd}$ and can no longer retain its magnetic flux (e.g., McClintock et al. 2006; Tchekhovskoy et al. 2014). This transition typically occurs hundreds of days post-disruption, consistent with the observed shut-off time of the X-rays in jetted TDEs (Zauderer et al., 2013; Cenko et al., 2012; Eftekhari et al., 2024). Super-Eddington accretion has also been proposed to power non-relativistic winds in TDEs that produce low-luminosity radio emission at early times (e.g., Strubbe and Quataert 2009; Alexander et al. 2016). The presence of a wind may explain why only powerful super-Eddington jets have been seen at early times: if the star’s orbital plane is initially misaligned with the black hole spin axis, the jet will precess and less powerful jets may only be able to break out of the wind after the disk undergoes sufficient Lense-Thirring precession to come into alignment with the spin (Teboul and Metzger, 2023; Lu et al., 2024). This will then delay any radio emission from less powerful super-Eddington jets. Radio emission from powerful relativistic jets may also appear delayed if the jets are viewed off-axis, as relativistic beaming will suppress the emission at early times. For nearby TDEs, off-axis jets may be directly resolved with Very Long Baseline Interferometry (VLBI) observations, providing a clear method for distinguishing between jetted and non-jetted emission (e.g., Mattila et al. 2018).
@@
Alternatively, a jet may form at very late times, once the accretion rate decreases to $\dot{M}\lesssim 0.03\,\dot{M}_{\rm Edd}$ (Giannios and Metzger, 2011). When the accretion rate becomes this low, the accretion flow is predicted to transition back to radiatively inefficient and geometrically thick, which can again power outflows (e.g., Narayan and Yi 1995; Blandford and Begelman 1999). Such behavior is supported by detections of radio jets from X-ray binaries (XRBs) in the low-hard state and bright radio emission consistent with compact jets from AGN at low Eddington ratios (e.g., Fender et al. 1999; Falcke et al. 2004; Sikora et al. 2007; Corbel et al. 2013). The rapid X-ray and optical spectral and photometric variability observed in changing-look AGN (Matt et al., 2003; Ricci and Trakhtenbrot, 2023) also supports the existence of this state change, as this variability is consistent with the sudden appearance or disappearance of a jet and changing-look AGN are predominantly observed to have Eddington ratios of a few percent, right at this expected transition. (This is much lower than the typical Eddington ratios seen in the general AGN population; MacLeod et al. 2019; Green et al. 2022; Guo et al. 2025.) A delayed radio jet was recently discovered in a changing-look AGN $\sim 6$ years after its original optical outburst, providing direct evidence that these events can power outflows (Meyer et al., 2025). In TDEs, this late jet phase is expected to occur on timescales of years to decades.
@@
We expect our TDEs to adhere to the Fundamental Plane of black hole activity (Section III.4) during periods where their SMBHs are accreting at a low fraction of the Eddington rate (Falcke et al., 2004; Jonker et al., 2020). Conversely, if the accretion rate is still high during some observed delayed radio brightenings, then we might expect the radio and X-ray emission to deviate from the Fundamental Plane for these TDEs. We may also expect to see a given TDE evolve along the Plane or towards/away from it as $\dot{M}$ changes and the accretion flow possibly undergoes state changes. The radio and X-ray emission of some TDEs also may not follow the Fundamental Plane at any epoch if the radio emission is not directly driven by accretion onto the SMBH (e.g., if prompt emission is produced instead by the shock between the unbound portion of the stellar debris stream and the ISM; Krolik et al. 2016, and/or if late-time emission is due to a previously-launched outflow encountering a change in the ambient environment; Zhuang et al. 2025). In the case of radio emission dominated by a precessing jet whose escape is delayed due to an initial misalignment between the disk and the SMBH spin, we also do not expect any clear correlation between the simultaneous radio and X-ray emission, as the radio onset time is determined by the alignment timescale rather than the accretion rate in these models (Teboul and Metzger, 2023; Lu et al., 2024).
IV.2 Origin(s) of the delayed radio emission
@@
Our delayed radio TDEs exhibit radio emission on a wide range of timescales (hundreds to thousands of days post-discovery) and with a wide range of luminosities, suggesting that a single mechanism may not be sufficient to explain the entire sample. In this section, we discuss the various types of accretion-driven outflows that may contribute to the delayed radio emission. As described in Section III, we constrain the accretion rate as a function of time for the TDEs in our sample via two methods: modeling the optical/UV light curves using MOSFiT, and comparing their X-ray and radio properties directly to the Fundamental Plane of black hole activity.
IV.2.1 (Delayed) super-Eddington accretion
@@
We first consider scenarios in which the delayed radio emission observed in our TDE sample is powered by non-relativistic outflows launched during a phase of super-Eddington accretion. While this is unlikely to explain TDEs with radio emission delayed by thousands of days (e.g., ASASSN-14ae) [^16] It is possible to delay peak accretion onto the SMBH by this much under certain assumptions (for example in the models for low-mass SMBHs explored by Guillochon and Ramirez-Ruiz 2015, where the debris stream may only self-intersect after dozens of pericenter passes and debris circularization can therefore be delayed by years). However, delaying the stream-stream collisions and debris circularization phases to this extent also spreads out the resulting mass fallback such that the peak accretion rate is likely no longer super-Eddington, and the optical evolution is likely much slower than observed in ASASSN-14ae., it may be a plausible explanation for TDEs where a radio outflow is launched hundreds days after discovery.
@@
Many of our delayed radio TDEs have larger optical photospheric radii at early times and delayed or absent He lines. As noted in Section II.3, previous work by Nicholl et al. (2022) found that TDE-H may arise from either less complete stellar disruptions (i.e., those with a smaller impact parameter, $b$) or from those in which an accretion disk does not form promptly. Our observed correlation between delayed radio emission and a TDE-H classification could thus suggest that less complete stellar disruptions are also more likely to produce delayed radio emission, but we disfavor this interpretation as we find no statistically-significant difference in the $b$ values of our Prompt and Delayed TDE sub-samples (Table 4). Instead, the fact that our delayed radio subsample both lacks prompt He emission and has larger optical photospheres favors the delayed accretion hypothesis. Nicholl et al. (2022) also found that their TDE-H exhibit larger photospheric radii than other types of TDEs, consistent with previous studies (Hinkle et al., 2020; van Velzen et al., 2021; Charalampopoulos et al., 2022). Nicholl et al. (2022) suggest that this could imply the early optical emission in TDE-H could be powered by collision-induced outflows rather than promptly formed accretion disks, such that accretion onto the SMHB is delayed relative to peak optical light. This would also naturally delay the onset of radio emission produced by accretion-powered outflows.
@@
Alternately, the stellar debris may circularize rapidly without immediately accreting. Instead, the circularized debris may initially form a hot quasi-spherical pressure-supported envelope (Steinberg and Stone, 2024; Price et al., 2024), delaying rapid accretion onto the SMBH until the envelope cools and contracts to form a disk (Loeb and Ulmer, 1997; Coughlin and Begelman, 2014; Metzger, 2022). If the pressure-supported envelope model is correct, the timing of the super-Eddington accretion phase may be delayed by several hundred days relative to the MOSFiT models presented in Section III.2, suggesting that it may still be possible for super-Eddington accretion to power the delayed radio emission observed in some TDEs even if the radio onset time occurs after MOSFiT’s nominal $\dot{M}\sim 0.3\;\dot{M_{\rm Edd}}$ transition point rather than prior to it (Metzger, 2022; Sarin and Metzger, 2024). In this case, the peak of the X-ray light curve will also be delayed relative to peak optical light and may provide a better estimate of the peak accretion time (Metzger, 2022). This is a likely explanation for the delayed X-rays seen in ASASSN-15oi (Figure 2 and Hajela et al. 2025). In scenarios where the early TDE UV/optical emission instead originates from a quasi-spherical envelope (Metzger, 2022), the initial size of the envelope (and hence its photosphere radius) increases with the SMBH mass; however, the cooling timescale over which the envelope contracts is also longer for lower SMBH mass, such that in the simplest scenario of passive cooling, the photosphere radius at a fixed time after the disruption is roughly independent of the SMBH mass. Given the low significance of our possible correlation between SMBH mass and delayed radio emission, this scenario is also consistent with our results.
IV.2.2 Relativistic jets
@@
The rapid radio brightening seen in these events may be powered by a jet (or wind) launched during a period of high accretion onto their SMBHs, when they are not expected to be on the Fundamental Plane. This is broadly consistent with the MOSFiT accretion rate predictions at the outflow launch times for ASASSN-15oi, PS16dtm, and AT2019ehz shown in Figures 2 and 3, although the systematic uncertainties on the SMBH masses for all three TDEs are large. The launch time for AT2018hyz shown in Figure 2 is taken from the non-relativistic outflow modeling by Cendes et al. (2022a) and does not coincide with MOSFiT’s prediction for super-Eddington accretion. This may either suggest that MOSFiT places the time of peak accretion too early, or that AT2018hyz’s radio emission is powered by an off-axis jet launched much earlier than the Cendes et al. (2022a) prediction but not visible until late times because of relativistic beaming (Matsumoto and Piran, 2023) or delayed disk-black hole spin alignment (Teboul and Metzger, 2023; Lu et al., 2024). Further radio monitoring of AT2018hyz, including VLBI observations, may distinguish among these possibilities.
@@
IV.2.3 Highly sub-Eddington accretion
@@
Some delayed radio emission in TDEs may also be powered by highly sub-Eddington accretion. To assess this possibility, we must accurately determine the timing of the $\dot{M}\;\sim 0.03\,\dot{M}_{\rm Edd}$ state transition, when the disk is expected to transition from geometrically thin back to geometrically thick. In our Case 1 prompt accretion scenario, we can straightforwardly compute this timescale using our $\dot{M}$ curves from MOSFiT (Section III.3). In Case 2 delayed accretion scenarios, the situation is less clear cut. If the cooling envelope model developed by Metzger (2022) is correct, then assuming that the envelope contracts to a disk faster than the mass fall-back reaches $\dot{M}\;\sim 0.03\,\dot{M}_{\rm Edd}$, and neglecting any spreading disk phase, the timescale estimates for this state transition that we obtain with MOSFiT while neglecting the envelope phase may still be accurate. [^18] We note that even if the MOSFiT prediction for the timing of the $\dot{M}\;\sim 0.03\,\dot{M}_{\rm Edd}$ transition is underestimated, parameters such as $R_{\rm ph,0}$ that are primarily determined by the photometry around peak optical light are still reliable. We find that some TDEs in our sample (e.g., AT2018hco) do indeed exhibit delayed radio emission consistent with an outflow launched at approximately the same time that MOSFiT predicts this state transition will occur (Figure 2).
@@
However, the existence of UV plateaus in many TDEs that remain well-fit by a thin disk model for years post-disruption (e.g., van Velzen et al. 2019b; Mummery and Balbus 2020; Mummery et al. 2025, 2024) may suggest that in reality the $\dot{M}\;\sim 0.03\,\dot{M}_{\rm Edd}$ transition is also delayed relative to the MOSFiT prediction, at least for some TDEs. [^19] This is not always the case; Mummery et al. (2025) apply the thin disk model FitTeD to AT2019dsg and find that the onset of its second radio flare approximately coincides with the calculated $\dot{M}\;\sim 0.03\,\dot{M}_{\rm Edd}$ state transition, just as MOSFiT does (Figure 3). Furthermore, in contrast to the hard X-ray spectra typical of AGN accreting at highly sub-Eddington rates, the X-ray emission from TDEs on timescales of years is often soft, consistent with thermal thin-disk emission extending to late times (e.g., Auchettl et al. 2017; Gezari et al. 2017; Wen et al. 2020; Jonker et al. 2020; Guolo et al. 2024 – but see also Komossa et al. 2004; Saxton et al. 2021 for discussion of X-ray selected TDEs that do show spectral hardening on timescales of years). We therefore must consider the full multiwavelength properties of each TDE in our sample to assess whether MOSFiT can provide an accurate estimate of its state transition timing, or whether other accretion probes (such as the X-ray emission) should be preferred.
@@
The strongest evidence that some of our delayed radio TDEs may have entered the highly sub-Eddington phase comes from comparing their radio and X-ray emission. Many of our delayed radio TDEs (e.g., ASASSN-14ae, AT2018fyk, and AT2019teq) align with the Fundamental Plane at late times (Figure 5). This may indicate that these TDEs have indeed transitioned to an analog of the low-hard state seen in XRBs at late times and, like the AGN and XRBs used to calibrate the original Fundamental Plane relationship, are accreting at a small fraction of the Eddington rate at the time of observation. Indeed, AT2018fyk’s and AT2019teq’s X-ray spectra show significant hardening at late times, consistent with this picture (Wevers et al., 2019a, 2021; Yao and Guolo, 2022; Guolo et al., 2024; Pasham et al., 2024). This is also consistent with the radio non-detections of AT2018fyk at very early times when the X-rays were significantly brighter and softer, as outflows are not expected prior to this state transition when the accretion rate is higher but likely not super-Eddington given the high SMBH mass (Fender et al. 1999; Fender and Gallo 2014; Wevers et al. 2019a). ASASSN-14ae is the only TDE in our sample to show a correlated increase in both radio and X-ray emission at late times (Figure 2), but its late-time X-ray emission remains too faint for a detailed spectral analysis (Jonker et al., 2020). ASASSN-14ae is also the oldest TDE in our sample and has the longest delay between peak optical light and first radio detection. As illustrated by Figure 2, in order to explain ASASSN-14ae’s radio emission as an outflow launched by a state transition in the accretion flow at $\dot{M}\sim 0.03\,\dot{M}_{\rm Edd}$, accretion onto the SMBH must be significantly delayed relative to the $\dot{M}$ curve computed by MOSFiT (by $\sim 2000$ days). Thus, further monitoring of ASASSN-14ae’s radio and X-ray emission is necessary to ascertain the nature of the late radio rise in this TDE.